Solve the system by the addition method: 2x=5 - 4y 4x - 4y=4. What is x?

Answer by ewatrrr(24785)

 
Hi
2x = 5 - 4y
4x - 4y=4
2x + 4y=5
4x - 4y=4
6x = 9
x = 9/6 = 3/2 and y = 1/2 y = (-2x+5)/4 = (-3+5)/4 = 1/2
CHECKING our Answer***
4x - 4y=4
6 - 2 = 4